DIRC Operations Report

Wednesday 2007/8/15 to Wednesday 2007/9/05


Oncall: Rolf Andreassen (Justine Serrano and Jose Benitez on backup)
Fast Monitoring: Rolf Andreassen, Justine Serrano, Jose Benitez

Contents:

    PEP Activities     BaBar Activities     DIRC activities     Backgrounds     Scheduled activities     Pending Work     Links

Luminosities this week

PEP peak lumi: ~1.17 x 1034 (last week: ~1.17 x 1034)

PEP delivery: 4.27 fb-1 (last week: 3.64 fb-1)

BaBar integrated lumi:3.90 fb-1 (last week: 3.47 fb-1)

Run 6 total delivery: 90.18 fb-1

Run 6 total BaBar recorded: 86.39 fb-1 total, including 7.89 fb-1 off-peak data

PEP activities/issues since last DIRC meeting

  • 08/16: Go back to on-resonance running.
  • 09/04: Shutdown for Run 6!

    BaBar activities/issues since last DIRC meeting

  • 08/27-29 : The narrower timing window for the SVT was introduced in an effort to reduce deadtime.

    DIRC activities/issues since last DIRC meeting

  • 08/15: Matt disabled the dump valve for the UPS battery replacement.
  • 08/17: Inhibited the chiller SIAM, which had been glitching again.
  • 08/22: As an experiment, tried to un-inhibit the chiller SIAM, since it seems to have been behaving nicely. It will get inhibited again if the glitches come back.
  • 08/25 onward: Some intermittently noisy channels making the shifters slightly nervous.
  • 08/27: Jochen and Chris tested the new FEX code with three colliding beam runs. Fast Mon and Prompt Calibration plots looked ok.
  • 08/29: Matt replaced the water filters.
  • 09/04: Took standalone runs to find Xmas trees.
  • 09/04: Shut off HV power supplies for shutdown.

    See DIRC logbook for details

    Lumi and Background since last DIRC meeting